Let L = {(1, 1, 1, 1, −4),(1, −1, 3, −2, −1)}. Find 6 vectors in the collection,
say H, such that L ∪ H spans the entire space.
1
Expert's answer
2019-09-26T11:12:35-0400

We can see that all vectors in "L\cup\{u_1,u_2,u_3,u_4,u_5,u_6,u_7,u_8\}" satisfy the equation "x_1+x_2+x_3+x_4+x_5=0".

Indeed, "l_1=(1,1,1,1,-4)" , so 1+1+1+1+(-4)=0

"l_2=(1,-1,3,-2,-1)" , so 1+(-1)+3+(-2)+(-1)=0

"u_1=(2,-3,4,-5,2)" , 2+(-3)+4+(-5)+2=0

"u_2=(-6,9,-12,15,-6)" , (-6)+9+(-12)+15+(-6)=0

"u_3=(3,\u22122,7,\u22129,1)" , 3+(-2)+7+(-9)+1=0

"u_4=(2,\u22128,2,\u22122,6)" , 2+(-8)+2+(-2)+6=0

"u_5=(\u22121,1,2,1,\u22123)" , (-1)+1+2+1+(-3)=0

"u_6=(0,\u22123,\u221218,9,12)" , 0+(-3)+(-18)+9+12=0

"u_7\u200b=(1,0,\u22122,3,\u22122)" , 1+0+(-2)+3+(-2)=0

"u_8\u200b=(2,\u22121,1,\u22129,7)" , 2+(-1)+1+(-9)+7=0

So all vectors in "L\cup\{u_1,u_2,u_3,u_4,u_5,u_6,u_7,u_8\}" satisfy the equation "x_1+x_2+x_3+x_4+x_5=0", that is "(1,1,1,1,1)\cdot (x_1,x_2,x_3,x_4,x_5)=0"

Let "T\vec{x}=(1,1,1,1,1)\cdot\vec{x}" , then "L\cup\{u_1,u_2,u_3,u_4,u_5,u_6,u_7,u_8\}\subset\ker T". Since "ImT\neq{0} (\bigl|(1,1,1,1,1)\bigr|^2\in ImT") and "ImT\subset R" , we have "\dim ImT=1"

Dimensional theorem (https://en.wikipedia.org/wiki/Dimension_theorem_for_vector_spaces) tells us that "\dim\ker T+\dim ImT=\dim V", that is "\dim\ker T+1=5", so "\dim\ker T=4".

Since "\langle L\cup\{u_1,u_2,u_3,u_4,u_5,u_6,u_7,u_8\}\rangle\subset\ker T", we have "\dim\langle L\cup\{u_1,u_2,u_3,u_4,u_5,u_6,u_7,u_8\}\rangle\le\dim\ker T=4", therefore "\langle L\cup\{u_1,u_2,u_3,u_4,u_5,u_6,u_7,u_8\}\rangle\neq V" . We cannot choose "H" such that "\langle L\cup H\rangle = V".